OpenStudy (awesomedude2078):

The account balance on April 1st is $60.15. On April 15th a payment of $51.00 is made. On April 25th a purchase of $91.27 is made. What is the finance charge if the annual rate is 18%? What is the new account balance? (Use the unpaid balance method.)

OpenStudy (awesomedude2078):

@AloneS

OpenStudy (awesomedude2078):

@mathmale

OpenStudy (wanell):

60.51 x 1.5% = 60.51(.015) = 0.91 <--finance charge 60.51 + 0.91 = 61.42 61.42 - 51.00 = 10.42 (subtract 51.00 payment 10.42 + 91.27 = 101.69 (add 91.27 purchase

OpenStudy (wanell):

is this right